OpenStudy (baby456):

simplify 4(3b+a-7) - (6b-7) help me and i will help you i prooooomise! Help MEEEEEEEEEEEEEEEEEEEEE! show steps and distribute

OpenStudy (luigi0210):

do you know how to distribute?

OpenStudy (baby456):

nope

OpenStudy (luigi0210):

well we have located the problem!

OpenStudy (zzr0ck3r):

then you need to go back and learn to distribute before you can do this...im guessing...

OpenStudy (zzr0ck3r):

a(b+c) = ab+ac <----- this is districution.

OpenStudy (zzr0ck3r):

distribution*

OpenStudy (baby456):

i know how just need help

OpenStudy (zzr0ck3r):

4(3b+a-7) distribute the 4 and what do you get?

OpenStudy (baby456):

12

OpenStudy (baby456):

12b

OpenStudy (zzr0ck3r):

12b+4a-28

OpenStudy (zzr0ck3r):

the 4 goes to all the terms inside()

OpenStudy (luigi0210):

what happened to your other variables? D:

OpenStudy (zzr0ck3r):

a(b+c) = ab+ac not ac

OpenStudy (baby456):

i need help

OpenStudy (zzr0ck3r):

4(3b+a-7) distribute the 4 and what do you get?

OpenStudy (baby456):

12b

OpenStudy (luigi0210):

no kidding, your variables and numbers just flew off into space :l

OpenStudy (zzr0ck3r):

I already told you that was wrong and then proceeded to tell you what the answer to that part was....you need to listen if you want help:)

OpenStudy (luigi0210):